# Artificial Intelligence HCI Human-Computer Interaction

### v1

@@ Abstract.
The advent of AI tools, such as Large Language Models, has introduced new possibilities for Qualitative Data Analysis (QDA), offering both opportunities and challenges. To help navigate the responsible integration of AI into QDA, we conducted semi-structured interviews with 15 HCI researchers experienced in QDA. While our participants were open to AI support in their QDA workflows, they expressed concerns about data privacy, autonomy, and the quality of AI outputs. In response, we developed a framework that spans from minimal to high AI involvement, providing tangible scenarios for integrating AI into HCI researchers’ QDA practices while addressing their needs and concerns. Aligned with real-life QDA workflows, we identify potentials for AI tools in areas such as data pre-processing, researcher onboarding, or mediation. Our framework aims to provoke further discussion on the development of AI-supported QDA and to help establish community standards for their responsible use.
@@
ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Collaborative interactionccs: Human-centered computing User studiesccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Recently, Artificial Intelligence (AI) tools have grown increasingly popular, exemplified by the rise of Large Language Models (LLMs). These powerful computational models can generate coherent, contextually relevant text across diverse topics and perform various tasks, including classification and summarization (Wei et al., 2022a; Bubeck et al., 2023). Large Language Models have found widespread adoption in numerous contexts, for example in workplace environments (e. g., assisting with programming or summarizing meeting notes) and education (e. g., helping explain complex concepts to students).
@@
Beyond these use cases, LLMs, alongside other AI tools, have the potential to transform Qualitative Data Analysis (QDA). Recent work has shown that Human-Computer Interaction (HCI) researchers already explore the usage of LLMs throughout their research processes for tasks such as conducting QDA (De Paoli, 2024), generating synthetic data and substituting human participants (Hämäläinen et al., 2023; Schmidt et al., 2024; Argyle et al., 2023), as well as for creative ideation (Shaer et al., 2024; Oppenlaender and Hämäläinen, 2023), literature reviews, and writing (Kapania et al., 2024). Some popular QDA software tools have already started incorporating AI features (e. g., nVivo’s Lumivero AI Assistant [^1] https://lumivero.com/product/nvivo-ai-assistant/(Morgan, 2023).
@@
The rapid advancement and adoption of AI tools by researchers beg the question: which tasks in the QDA process are (not) suitable for automation by AI tools, and how do we make these determinations? While AI tools are useful, they raise numerous dilemmas, including how AI might undermine the role of the researcher (Feuston and Brubaker, 2021), and how to deal with imperfect AI tools (e. g., LLMs hallucinating wrong information) (Xiao and Wang, 2021). As a community, it is imperative that we critically assess which tasks in the QDA workflow are appropriate or inappropriate for AI integration, and ensure their responsible use. To better understand how AI tools can, and should, be integrated into researchers QDA workflows, we conducted semi-structured interviews with HCI researchers to ask about their QDA workflows and opinions on AI tools. We find that while researchers’ QDA workflows share structural similarities, they vary significantly depending on project specifics, necessitating AI tools that are highly customizable. Moreover, HCI researchers are open to AI support if it enhances their workflow but emphasize the importance of ethical considerations, trust in AI outputs, and maintaining control over the analysis process. We then used the findings from these interviews to ideate tangible scenarios for integrating AIs into QDA in ways that support the QDA process while focusing on researchers’ needs, including AI as a productivity tool (e. g., for calculating inter-rater reliability), as a mediator (e. g., to help researchers reach consensus efficiently), or with human approval (e. g., where an AI autonomously creates and suggests codebook components, which are then approved or rejected by a researcher).
Our contributions are as follows. First, we offer deep insights into how researchers conduct QDA, highlighting their primary pain points and challenges. Next, we present a framework for incorporating AI into researchers’ workflows, contextualized with their perspectives on the individual stages of QDA. We discuss how this framework can be applied to responsibly deploy AI within QDA. Together, these provide insights and guidance for those developing AI tools for QDA, ensuring they are implemented in ways that are genuinely useful to researchers.
@@ 2. Background
In the following sections, we lay out the foundational backgrounds of QDA, and present recent developments in the space of AI-assisted QDA. We finish by synthesizing our research questions and contributions.
2.1. Qualitative Data Analysis
@@
QDA is a fundamental process in qualitative and mixed-method research, encompassing techniques for systematic extraction and interpretation of non-numeric data, such as interview transcripts, field notes, videos, or textual documents (Booth et al., 2012). Its primary goal is to juxtapose or integrate different sources to identify patterns, themes, and insights that contribute to a deeper understanding of the research subject. At its core, QDA typically involves three steps that can also be applied iteratively: data extraction, synthesis, and analysis (Booth et al., 2012).
@@
QDA is shaped by various methodological frameworks that may overlap in scientific practice (see reporting of QDA steps in Section 4.1) but are distinct in theoretical foundations, flexibility, and depth of analysis. While Thematic Analysis (Braun and Clarke, 2006) is a theoretical framework that focuses on a coding process in which researchers highlight features of the data relevant to the research question, Content Analysis (Wilkinson, 2000) quantifies specific events or codes, enabling a blend with quantitative approaches. Grounded Theory (Glaser and Strauss, 1967) aims at theory generation through systematic coding directly from the data, and Framework Analysis (cf. Barnett-Page and Thomas, 2009) applies existing frameworks, refining them during analysis where needed, and is frequently employed in applied research contexts that aim to build upon and extend existing knowledge. In contrast, Narrative Analysis (Riessman, 2008) emphasizes the structure and meaning of data within its narrative context.
On a higher level, coding strategies can be either deductive (fitting data to pre-existing categories) or inductive (allowing patterns to emerge from the data). Some methods accommodate both strategies, enabling researchers to choose an approach that aligns with their objectives or combine elements of both for a more comprehensive analysis. The choice of coding strategy is not merely a technical or skill-related decision but reflects the researcher’s underlying goals and questions, guiding the research direction and ensuring the chosen method supports the intended outcomes.
A significant research gap exists in understanding how these distinctions are navigated in practice, lacking empirical studies examining HCI researchers’ choices and implementations of these theoretical methods. This lack of clarity limits methodological innovation and the development of best practices, particularly regarding the integration of computer or AI assistance, highlighting the need for more systematic investigation into current QDA practices.
2.2. Computer-Assisted QDA: From Basic NLP Methods to LLM-Driven Qualitative Analysis
Computer-assisted approaches are increasingly integrated into QDA, beginning with early methods using basic Natural Language Processing (NLP) techniques like word frequency analysis, keyword matching, and pattern-based rules (Stone et al., 1962; Marathe and Toyama, 2018). While these methods introduced automation, they lacked the capacity to capture complex human interpretations. The development of supervised machine learning models advanced computer-assisted QDA by enabling more sophisticated text classification based on labeled data. For instance, Rietz et al. (Rietz and Maedche, 2021) developed an AI-based system that allows researchers to define rules and use supervised learning to extend manual coding to unseen data. However, these models require large amounts of annotated data, which can be difficult to obtain, particularly for rare codes. Unsupervised machine learning methods classify text based on inherent features and patterns and do not require labeled data. Lennon et al. (Lennon et al., 2021), for example, utilized an unsupervised, graph-based approach to model topic categories. However, despite these models’ ability to discover themes without predefined categories, they often lack transparency, raising concerns about the interpretability of their coding decisions in QDA contexts. This limitation is particularly problematic in qualitative research, where the goal is precisely that: extract insights and achieve a deep understanding of the data’s underlying patterns and context.
Recent advances in generative AI, particularly LLMs like ChatGPT, offer new opportunities for QDA by providing tools for text generation and analysis (Wei et al., 2022b). Trained on large and diverse corpora of literature, these models encode vast amounts of knowledge and can understand context in nuanced ways, potentially making them adaptable to various tasks within QDA. LLMs have shown remarkable performance in annotation tasks where little or no labeled data is given (Gilardi et al., 2023; Törnberg, 2023). Ziems et al. (Ziems et al., 2024) highlight the potential for LLMs to serve as data annotators in human annotation teams. The integration of AI features into QDA software such as ATLAS.ti [^2] https://atlasti.com/ and MAXQDA [^3] https://www.maxqda.com/products/ai-assist also reflects the growing interest in leveraging AI to facilitate (semi)-automated coding processes. A range of emerging studies and prototypes now explore incorporating LLMs at different stages of the QDA workflow, with varying levels of human interaction and oversight, suggesting a need for tailored approaches that align with the specific needs of qualitative researchers (Bano et al., 2024; Tai et al., 2024; Xiao et al., 2023; Morgan, 2023; Rasheed et al., 2024).
2.3. AI in the QDA Workflow
@@
Recent studies have explored various ways AI can enhance QDA, particularly through applications like automated labeling, code prediction, and codebook development. For labeling and annotating, AI systems have shown promise in applying codes from an existing codebook consistently, reducing the time needed for manual coding. For example, Xiao et al. (Xiao et al., 2023) demonstrated that models like GPT-3 can achieve fair to substantial agreement with expert coders, although they may struggle with more nuanced human judgments and contextual understanding. For data exploration and inductive coding, LLMs can aid in the initial exploration of themes and codes by identifying novel patterns within the data, potentially enabling a deeper understanding. While LLMs can replicate many primary themes identified by humans, they often miss more subtle and complex interpretations (Morgan, 2023; Zhang et al., 2023; Torii et al., 2024). Hamilton et al. (Hamilton et al., 2023) also observed differences between human and ChatGPT coding, with each identifying themes that the other did not. Barany et al. (Barany et al., 2024) investigated the use of ChatGPT for codebook creation, finding hybrid approaches to be more reliable. They suggest LLMs can contribute to the improvement of codebooks, but stress that human involvement remains crucial. Furthermore, AI can also play a role in facilitating collaboration among researchers by mediating discussions, highlighting coding discrepancies, and suggesting resolutions (Gao et al., 2023; Ganji et al., 2018; Drouhard et al., 2017).
2.4. Design for Human-AI Collaboration
@@
Human-AI collaboration is gaining popularity across various fields as AI tools are developed to streamline repetitive tasks and enhance decision-making. QDA presents unique challenges for AI integration due to its interpretive and context-dependent nature (Jiang et al., 2021). Existing tools for QDA vary in their levels of AI support, from minimal assistance to advanced roles in coding and analysis, with different levels of human involvement. While prior studies provide valuable insights into the potentials of AI in QDA (Tai et al., 2024; Morgan, 2023; Rasheed et al., 2024), there is still limited understanding of how AI can be responsibly integrated into researchers’ real-life QDA workflows, respecting their concerns and preferences (Lubars and Tan, 2019; Feuston and Brubaker, 2021). To address these gaps, we study HCI researchers’ QDA workflows and their perspectives on the involvement of AI in qualitative research. Specifically, we investigate the following research questions:
@@
Our overarching goal is to explore how AI can be responsibly integrated into QDA, respecting the workflows, concerns, and conditions of HCI researchers. Based on our detailed insights from RQ1 and RQ2 and previous work, we synthesize a framework for AI-based support that is aligned with the real-world QDA workflows of HCI researchers. We anticipate that our framework can contribute to the future design and evaluation of AI tools that enhance – rather than disrupt – qualitative analysis.
@@ 3. Research Method
We conducted a two-step study to explore HCI researchers’ perspectives on integrating AI into QDA. The first step involved semi-structured interviews with 15 HCI researchers, representing a range of QDA experience levels. These interviews were designed to uncover pain points, potentials, and specific conditions for AI support in QDA, focusing on real-life workflows, challenges, and concerns. To enhance the transparency and rigor of our research process, we have pre-registered our study on the Open Science Framework (OSF) platform.444see https://osf.io/yzkta/?view_only=c4bb80ff8635434db0e7c3166146fa53 For the second step, we used the findings of the interviews to conceptualize ways AI can be integrated into QDA. We developed a framework to illustrate tangible, concrete examples of how AI can be integrated into the QDA workflow and contextualize these examples with how they align (or do not align) with participants’ concerns. We now describe our interview method in detail.

### v2

@@ Abstract.
The advent of Artificial Intelligence (AI) tools, such as Large Language Models, has introduced new possibilities for Qualitative Data Analysis (QDA), offering both opportunities and challenges. To help navigate the responsible integration of AI into QDA, we conducted semi-structured interviews with 15 Human-Computer Interaction (HCI) researchers experienced in QDA. While our participants were open to AI support in their QDA workflows, they expressed concerns about data privacy, autonomy, and the quality of AI outputs. In response, we developed a framework that spans from minimal to high AI involvement, providing tangible scenarios for integrating AI into QDA practices while addressing researchers’ needs and concerns. Aligned with real-life QDA workflows, we identify potential for AI tools in areas such as data pre-processing, researcher onboarding, or conflict mediation. Our framework aims to provoke further discussion on the development of AI-supported QDA and to help establish community standards for responsible Human-AI collaboration.
@@
$c$s: Human-centered computing Interactive systems and toolsccs: Human-centered computing Collaborative and social computing design and evaluation methodsccs: Human-centered computing Collaborative interactionccs: Human-centered computing User studiesccs: Human-centered computing HCI theory, concepts and modelsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Recently, Artificial Intelligence (AI) tools have grown increasingly popular, driven by the rapid rise of Large Language Models (LLMs). These powerful models can generate coherent, contextually relevant text across diverse topics and perform tasks such as classification and summarization (Wei et al., 2022a; Bubeck et al., 2023). Large Language Models have been widely adopted across domains, including workspaces, from programming assistance to educational tutoring, and are increasingly finding their way into research contexts.
@@
One area where these models hold significant potential is Qualitative Data Analysis (QDA). Human-Computer Interaction (HCI) researchers are beginning to explore the usage of LLMs throughout their research processes for tasks such as conducting QDA (De Paoli, 2024), generating synthetic data and substituting human participants (Hämäläinen et al., 2023; Schmidt et al., 2024; Argyle et al., 2023), creative ideation (Shaer et al., 2024; Oppenlaender and Hämäläinen, 2023), literature reviews, and writing (Kapania et al., 2024a). Commercial tools have also started integrating AI features (e. g., ATLAS.ti, [^1] https://atlasti.com/ MAXQDA, [^2] https://www.maxqda.com/products/ai-assist and nVivo’s Lumivero AI Assistant [^3] https://lumivero.com/product/nvivo-ai-assistant/(Morgan, 2023).
@@
The rapid adoption of AI tools begs the question: which parts of the QDA workflow are suitable for AI support, and how do we make these determinations? These questions are particularly pivotal for CSCW, where QDA plays a dual role: as a methodological tool, and as a site of collaboration, negotiation, and interpretive practice. Although AI tools can be useful, they raise dilemmas about researcher autonomy (Feuston and Brubaker, 2021) and the risks of imperfect AI outputs (e. g., LLMs hallucinating wrong information) (Xiao and Wang, 2021). As a community, it is imperative that we critically assess which tasks in the QDA workflow are appropriate or inappropriate for AI integration, and to ensure its responsible use. To better understand how AI tools can, and should, be integrated into researchers’ QDA workflows, we conducted semi-structured interviews with HCI researchers about their QDA workflows and opinions on AI integration. We found that while researchers’ QDA workflows share structural similarities, they differ significantly based on project specifics, highlighting the need for AI tools that are highly customizable. Researchers were generally open to AI assistance when it enhanced their workflow but stressed the importance of ethical considerations, trust in AI outputs, and maintaining control.
Our contributions are as follows. First, we provide an empirical account of researchers’ QDA practices, pain points, and challenges. Second, we present a framework for integrating AI into QDA workflows, grounded in researchers’ perspectives on the individual stages of QDA. Based on our findings we ideate tangible scenarios for the integaration of AI, including AI as a productivity tool (e. g., calculating inter-rater reliability), as a mediator (e. g., facilitating consensus), or with human approval (e. g., suggesting codebook components with validation by a researcher). Finally, we discuss how this framework can guide the responsible deployment of AI tools that truly support researchers’ needs and preserve the social value of collaborative analysis practices.
@@ 2. Background
In this section, we lay out the methodological backgrounds of Qualitative Data Analysis (QDA) and trace the evolution of AI-assisted QDA from basic computational methods to the growing role of Large Language Models (LLMs). We conclude by synthesizing our research questions.
2.1. Qualitative Data Analysis in CSCW and HCI
@@
QDA is a fundamental process in qualitative and mixed-method research within CSCW and HCI, providing systematic techniques for interpreting non-numeric data such as interviews, field notes, videos, or documents (Booth et al., 2012; McDonald et al., 2019). Its primary goal is to juxtapose or integrate different sources to identify patterns, themes, and insights that deepen our understanding of the research subject. At its core, QDA typically involves iterative cycles of data extraction, synthesis, and analysis (Booth et al., 2012).
@@
QDA is shaped by various methodological frameworks that may overlap in scientific practice (see reporting of QDA steps in Section 4.1) but are distinct in theoretical foundations, flexibility, and depth of analysis. While Thematic Analysis (Braun and Clarke, 2006) is a theoretical framework that focuses on coding features relevant to the research question, Content Analysis (Wilkinson, 2000) quantifies specific events or codes, enabling a blend with quantitative approaches. Grounded Theory (Glaser and Strauss, 1967) aims at theory generation through systematic coding directly from the data, and Framework Analysis (cf. Barnett-Page and Thomas, 2009) applies existing frameworks, refining them throughout analysis, useful in applied research contexts that aim to build upon and extend existing knowledge. In contrast, Narrative Analysis (Riessman, 2008) emphasizes the structure and meaning of data within its narrative context. Coding strategies can be deductive (codes follow predefined categories), inductive (patterns identified in the data), or hybrid, enabling researchers to choose an approach that aligns with their objectives. The choice of coding strategy is not merely technical but reflects a researcher’s underlying goals and questions, guiding the research direction and ensuring the chosen method supports the intended outcomes.
Beyond methods, QDA in CSCW and HCI is shaped by interdisciplinary diversity. Researchers bring different epistemological traditions and backgrounds (Fiesler et al., 2019; McDonald et al., 2019), which leads to ongoing debates about transparency, rigor, reliability, and measures thereof, such as Inter-Rater Reliability (IRR),  (Fiesler et al., 2019; McDonald et al., 2019). At the same time, most CSCW and HCI projects involve more than one researcher (McDonald et al., 2019), making QDA a fundamentally collaborative and social practice of negotiating, resolving disagreements, and making sense of ambiguity. This collaboration is often asynchronous and distributed, and supported by digital tools and software (Wallace et al., 2017), increasingly incorporating Natural Language Processing (NLP) and LLM methods. Yet, despite QDA’s centrality, there has been no systematic documentation of how CSCW and HCI researchers actually conduct QDA in practice, or how emerging technologies might reshape it. Our work addresses this gap by synthesizing and extending fragmented conversations into a coherent account of real-world QDA workflows.
2.2. From Computational Support to AI-Assisted Analysis
While CSCW and HCI researchers draw on a variety of tools for analysis (Fiesler et al., 2019) computer-assisted approaches have long supported QDA. Early tools applied basic NLP techniques such as word frequency analysis, keyword matching, and pattern-based rules (Stone et al., 1962; Marathe and Toyama, 2018). While these methods introduced automation, they lacked the capacity to capture complex human interpretations. Subsequent work introduced supervised machine learning for text classification, extending manual coding into larger datasets (Rietz and Maedche, 2021). Unsupervised approaches such as topic modeling uncovered latent themes without predefined categories (Lennon et al., 2021), and some studies reported surprising alignment with qualitative methods like grounded theory (Baumer et al., 2017; Muller et al., 2016). Such combinations of QDA and machine learning can be especially useful for mixed-methods studies and for analyzing large-scale data, though they raise concerns about transparency and interpretability – values central to qualitative research. Recent advances in generative AI, particularly LLMs like ChatGPT, offer new opportunities. These models are trained on massive text corpora and can generate, summarize, and classify language with a high degree of contextual awareness (Wei et al., 2022b). They encode vast amounts of diverse knowledge, making them potentially adaptable to various tasks within qualitative research. LLMs have shown remarkable performance on annotation tasks where little or no labeled data is provided (Gilardi et al., 2023; Törnberg, 2023a) and have been proposed as collaborators in human annotation teams (Ziems et al., 2024a). Recent prototypes integrate LLMs at different points across the QDA workflow, from inductive coding to codebook development (Bano et al., 2024; Tai et al., 2024; Xiao et al., 2023; Morgan, 2023; Rasheed et al., 2024; Gauthier and Wallace, 2022). This evolution highlights not only the opportunities for scaling qualitative work but also the risks it creates for preserving interpretive rigor.
2.3. Potentials and Risks of AI in the QDA Workflow
@@
Recent studies show how AI can assist at multiple points in the QDA process. Cuevas et al. (2025) examine the usage of LLMs for conducting user studies and analyzing the collected data. For labeling and annotating, LLM-based systems apply existing codes from a given codebook with notable consistency, though they often miss subtle contextual judgments (Xiao et al., 2023). Lam et al. (2024) develop a LLM-based text analysis tool that derives interpretable concepts from unstructured text, inspired by traditional topic models. For data exploration and inductive coding, LLMs can surface novel patterns and assist in the initial exploration of themes and codes, potentially enabling a deeper understanding. However, while LLMs can replicate many primary themes identified by humans, they often miss more subtle and complex interpretations (Morgan, 2023; Zhang et al., 2023; Torii et al., 2024). Hamilton et al. (2023) observed complementary patterns between human and ChatGPT-generated codes, with each identifying themes that the other did not. Barany et al. (2024) investigated the use of ChatGPT for codebook creation, finding hybrid approaches to be more reliable. They suggest LLMs can contribute to the improvement of codebooks, but stress that human involvement remains crucial. While combining interpretive and positivist methods can cause epistemic tensions and discussions among researchers (Baumer et al., 2017; Muller et al., 2016; Paulus and Marone, 2024), AI also holds potential for mediating discussions and facilitating collaboration by highlighting coding discrepancies, and suggesting resolutions (Gao et al., 2023; Ganji et al., 2018; Drouhard et al., 2017).
While these efforts illustrate potential, they also expose tensions. Concerns persist around bias (Bail, 2024; Törnberg, 2023b; Ziems et al., 2024b; Anis and French, 2023; Owoahene Acheampong and Nyaaba, 2024; Bano et al., 2023), lack of reliability for non-English languages (Heseltine and Clemm von Hohenberg, 2024; Suter and Meckel, 2024), and the risk of treating AI as “culturally agnostic” despite its privileging of certain perspectives (Paulus and Marone, 2024). Such issues are particularly important when engaging Global South contexts, where AI risks reproducing inequalities in representation (Kumar and Karusala, 2021; Wacharamanotham et al., 2020; Rifat et al., 2023; Jang et al., 2023; Okolo et al., 2022). Thus, while AI holds potential to bridge “big” and “small” data approaches and to enhance efficiency, it raises equally urgent questions about reliability, ethics, and researcher agency.
2.4. Designing for Human-AI Collaboration
Integrating AI into QDA is not simply a matter of automation. It requires careful design for human–AI collaboration in interpretive, context-sensitive work (Jiang et al., 2021). Existing tools vary widely in how they support AI involvement, from minimal assistance to active roles in coding and analysis, with different levels of human involvement. Moreover, designing for researchers differing in proficiency in tool usage as well as AI capability (Lubars and Tan, 2019; Feuston and Brubaker, 2021) poses specific challenges, especially in diverse fields such as CSCW and HCI. For example, Gauthier and Wallace (2022) developed a toolkit to facilitate computational thematic analysis, emphasizing the need to design it accessibly for a community with different proficiencies in tool usage as well as AI. Thus, they call for ethical designs to involve friction to remind researchers of the imperfection of computational models and issues concerning ethics and transparency. These challenges are also identified by Kapania et al. (2024b) who found HCI researchers struggling to identify such issues in their projects and needing guidance. Similarly, Lam et al. (2024) suggest tools to be transparent to the researcher and acknowledging their limitations, e. g., by showcasing explicit inclusion criteria and scores. At the same time, designs should avoid creating excessive “unanticipated labor” for researchers (Feuston and Brubaker, 2021).
@@ 2.4. Designing for Human-AI Collaboration
While prior studies provide valuable insights into the potentials of AI in QDA (Tai et al., 2024; Morgan, 2023; Rasheed et al., 2024; Lam et al., 2024; Gauthier and Wallace, 2022), there is still limited understanding of how AI can be responsibly integrated into researchers’ real-life QDA workflows, respecting their concerns and preferences (Lubars and Tan, 2019; Feuston and Brubaker, 2021). We are thus in need for comprehensive frameworks addressing these issues (Owoahene Acheampong and Nyaaba, 2024; Kapania et al., 2024b). To address these gaps, we study HCI researchers’ QDA workflows and their perspectives on the involvement of AI in qualitative research. Specifically, we investigate the following research questions:
@@
Our overarching goal is to explore how AI can be responsibly integrated into QDA, respecting the workflows, concerns, and conditions of HCI researchers. By systematically mapping QDA practices (RQ1) and surfacing researchers’ conditions for AI integration (RQ2), we contribute a framework for AI support that is aligned with the real-world QDA workflows of HCI researchers. This framework synthesizes fragmented CSCW debates around QDA and provides concrete guidance for developing AI tools that enhance – rather than disrupt – the collaborative and interpretive nature of qualitative research.
@@ 3. Research Method
We conducted a two-step study to explore HCI researchers’ perspectives on integrating AI into QDA. The first step involved semi-structured interviews with 15 HCI researchers, representing a range of QDA experience levels. These interviews were designed to uncover pain points, potentials, and conditions for AI support in QDA, focusing on real-life workflows, challenges, and concerns. To promote transparency and rigor, we preregistered our study on the Open Science Framework (OSF) platform.444see https://osf.io/yzkta/?view_only=c4bb80ff8635434db0e7c3166146fa53 In the second step, we used the findings of the interviews to conceptualize ways AI can be integrated into QDA. We developed a framework to illustrate concrete examples of how AI can be integrated into the QDA workflow and contextualize these examples with how they align (or do not align) with participants’ concerns.